OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the given point. y = 5x − 2squarerootx, (1, 3)

OpenStudy (mertsj):

Can you find the first derivative of the function?

OpenStudy (mertsj):

@ElfQueen Hello. Are you there?

OpenStudy (anonymous):

You may find the first derivative and then find f'(1) which will give you your slope, and then use the following equation \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

@Mertsj Sorry. I was making myself dinner. I got 5 - 1over the squareroot x.

OpenStudy (anonymous):

That's correct \[f'(x) = 5-\frac{ 1 }{ \sqrt{x} }\] now find f'(1)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes! So our slope is 4, now we can use the point slope equation and we will have our equation

OpenStudy (anonymous):

4x-1

OpenStudy (anonymous):

I'm not sure what that means, we just have to use \[y-y_1=m(x-x_1)\] and we are given the point (x1,y1)